OpenStudy (anonymous):

eliminate the arbitrary constant y=c1x^2 + c2e^-x c1= c sub 1 c2= c sub 2 ty

OpenStudy (dan815):

??

OpenStudy (dan815):

anymore info?

OpenStudy (dan815):

set c1 and c2 = some random number?

OpenStudy (dan815):

y=0 is a solution

OpenStudy (perl):

take the derivatives

OpenStudy (perl):

y=c1x^2 + c2e^-x dy/dx = c1*2x - c2*e^(-x) d^2y/dx^2 = c1*2 + c2*e^(-x) d^3 y/dx^2 = - c2*e^(-x)

OpenStudy (perl):

now we can eliminate c1 and c2

OpenStudy (perl):

y=c1x^2 + c2e^-x y ' = c1*2x - c2*e^(-x) y ' ' = c1*2 + c2*e^(-x) y ' ' ' = - c2*e^(-x)

OpenStudy (perl):

notice that y ' ' + y ' ' ' = 2*c1

OpenStudy (anonymous):

yeah, i got the whole thing, thank you so much master perl

OpenStudy (perl):

:D

OpenStudy (anonymous):

just to confirm, the derivative of 2c1 is 2 and not c1?

OpenStudy (perl):

no, the derivative of 2*c1 is zero, since it is a constant times constant

OpenStudy (perl):

i made the same mistake, i thought it represented a variable. but c sub 1 and c sub 2 represent arbitrary constants (which you pick)

OpenStudy (perl):

a constant times a constant will be a constant, so the derivative of 2 * c_1 is 0

OpenStudy (perl):

y=c1x^2 + c2e^-x y ' = c1*2x - c2*e^(-x) y ' ' = c1*2 + c2*e^(-x) y ' ' ' = 0 - c2*e^(-x)

OpenStudy (perl):

wait, woops

OpenStudy (anonymous):

yeah, it's so clear to me now, i can handle it from here, again thank you so much master perl

OpenStudy (perl):

we are looking for an ODE that has no arbitrary constants c_1 , c_2

OpenStudy (perl):

which satisfies y = C_1*x^2 + C_2 * exp(-x)

OpenStudy (perl):

y=c1x^2 + c2e^-x y ' = c1*2x - c2*e^(-x) y ' ' = c1*2 + c2*e^(-x) y ' ' ' = - c2*e^(-x) y '''' = c2*e^(-x). therefore y ' ' ' + y ' ' ' ' = 0

OpenStudy (perl):

and according to wolfram the solution of y ' ' ' + y ' ' ' ' = 0 is equal to y(x) = c_4 x^2+c_3 x+c_1 e^(-x)+c_2 Now set c_3 = 0 and c_2 = 0, and we have our solution y given above. http://www.wolframalpha.com/input/?i=solve+y%27%27%27+%2B+y%27%27%27%27+%3D+0+

OpenStudy (perl):

@dan815 want to check this makes sense?

OpenStudy (anonymous):

how can i thank you more? you're just so awesome, by any chance are you a prof?

OpenStudy (perl):

again, to be clear what the question is asking Find a differential equation that y satisfies. The ODE should not contain any arbitrary constants. ANd y = c_1*x^2 + c_2 * e^(-x)

OpenStudy (perl):

not yet a professor

OpenStudy (anonymous):

yeah, been trying to eliminate that for about 4 hrs now and here you are, you saved me, you're gonna be a great prof perl

OpenStudy (perl):

thanks :D

OpenStudy (dan815):

nice :)

OpenStudy (perl):

but next midorima, give more information in the question !!! then you might get more replies

OpenStudy (perl):

next time

OpenStudy (anonymous):

sure i will :)

OpenStudy (perl):

if you have any questions about this one, let me know